OpenStudy (anonymous):

how do i do this???? write as a single fraction -1 +4c-9d/6c - 5c-4d/8c

OpenStudy (anonymous):

/ = divided by

OpenStudy (tkhunny):

It would help to clarify exactly what you have written. What I see is this: \(-1 + 4c - \dfrac{9d}{6c} - 5c - \dfrac{4d}{8c}\) If this is not what you intended, please use more parentheses to clarify intent. Remember your Order of Operations.

OpenStudy (anonymous):

-1 + (4c-9d/6c) (5c-4d/8c)

OpenStudy (anonymous):

oops I meant minus between the two parenthesis not multiply

OpenStudy (anonymous):

-1 + (4c-9d/6c) - (5c-4d/8c)

OpenStudy (tkhunny):

This, then: \(-1 + \left(4c - \dfrac{9d}{6c}\right) - \left(5c - \dfrac{4d}{8c}\right)\)

OpenStudy (anonymous):

no 4c-9d is the top of the fraction and 6c is the bottom of the fraction

OpenStudy (anonymous):

5c-4d top of fraction and 8c bottom of fraction

OpenStudy (tkhunny):

You have forgotten your Order of Operations, haven't you? -1 + (4c-9d)/6c - (5c-4d)/8c = \(-1 + \dfrac{4c-9d}{6c} - \dfrac{5c-4d}{8c}\) Are we close, yet?

OpenStudy (anonymous):

yes! that is it

OpenStudy (tkhunny):

What is the least common denominator?

OpenStudy (anonymous):

1?

OpenStudy (tkhunny):

Back up. When adding these, how might one proceed? \(\dfrac{1}{8} + \dfrac{1}{6}\) Common denominator?

OpenStudy (anonymous):

24?

OpenStudy (tkhunny):

Okay, how about these: \(1 + \dfrac{1}{3} + \dfrac{1}{4}\)

OpenStudy (anonymous):

12?

OpenStudy (tkhunny):

Good. Please add those three for me.

OpenStudy (anonymous):

3/24?

OpenStudy (tkhunny):

?? How did you get that? \(1 + \dfrac{1}{3} + \dfrac{1}{4} = \dfrac{12}{12} + \dfrac{4}{12} + \dfrac{3}{12} = \dfrac{19}{12}\) This is nowhere near what you just said. You need to be able to add fractions. Please show me \(\dfrac{1}{8} + \dfrac{1}{6}\). Add them.

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

all I want is the answer, im confused. and I don't understand fractions and never will

OpenStudy (tkhunny):

Therein lies the problem. Why do you have an algebra problem to add algebraic fractions when it appears you have no background in adding numerical fractions? 1) Find a common denominator. 2) Promote each term to the common denominator. 3) Sum the numerators over the common denominator. \(\dfrac{1}{8} + \dfrac{1}{6}\) Common Denominator: 24 \(\dfrac{1}{8} + \dfrac{1}{6} = \dfrac{3}{24} + \dfrac{4}{24}\) Fractions promoted to the common denominator. \(\dfrac{3+4}{24} = \dfrac{7}{24}\) You need to be able to do this fluently. Numerators summed and it is still over that same denominator.

OpenStudy (tkhunny):

If you do not understand fractions and never will, then you cannot do this problem. That is the best advice anyone can give you.